OpenStudy (anonymous):

which equation has a solution of -6? F. 15x-20=70 G. 14=6x-22 H. 3x-8=-10 I. 1 over 2x-8=-11

OpenStudy (anonymous):

helpp plzzz

OpenStudy (anonymous):

15x=90

OpenStudy (anonymous):

x=90/15

OpenStudy (anonymous):

x=6

OpenStudy (anonymous):

F

OpenStudy (anonymous):

oh wait -6

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

I

OpenStudy (anonymous):

the answer is I

OpenStudy (anonymous):

1/2x=-3

OpenStudy (anonymous):

2(1/2x=-3)

OpenStudy (anonymous):

x=-6